# Jouko Väänänen Helsinki and Amsterdam Ur Ya’ar

### v2

@@ 3 The quantifier $Q_{1}$
As a second example, let $Q_{1}$ be the quantifier asserting “there are $\geq\aleph_{1}$ many…”. It is well-known that $\Delta(\mathcal{L}(Q_{1}))\neq\mathcal{L}(Q_{1})$, as demonstrated by the property of an equivalence relation having uncountably many uncountable classes (cf. [BF85, Chapter II section 7.2]). This raises the question, what is the relationship between $C(\mathcal{L}(Q_{1}))$ and $C(\Delta(\mathcal{L}(Q_{1})))$?
@@
The proof will use the following machinery. By [AS93, Theorem 7.3], there is a sequence of Suslin trees $\mathcal{T}=\left\langle T_{n}\mid n<\omega\right\rangle\in L$ such that for every $i\neq j$, $T_{i}\otimes T_{j}$ is Suslin. It is well known (see [Jec07, Lemma 16.5]) that if $S,T$ are Suslin trees such that $S\otimes T$ is Suslin as well, then forcing with one (to add a generic branch) preserves the Suslinity of the other. This then implies, by [AS93, Theorem 4.6], that there is a proper forcing notion which specializes one while keeping the other Suslin. So, all in all, we can either force a branch or specialize $T_{i}$ without affecting any of the other $T_{j}$s.
@@
Let $a\subset\omega$ be a Cohen real over $L$. Using a finite support product, we force over $L[a]$ to add a branch to any $T_{i}$ for $i\in a$ and specialize any $T_{j}$ for $j\notin a$. Then we force Martin’s Axiom ($\mathsf{MA}$) in the standard way over the resulting model, and denote the final model $W$. Note that $\omega_{1}^{W}=\omega_{1}^{L}$. We now work in $W$. As a consequence of $\mathsf{MA}$, any Aronszajn tree in $W$ is special, and we have that
@@ 6 Conclusion and open questions
  4. 4.
    Is it consistent that $L\neq V=C(\Delta(Q_{1}))(\text{or }C(\Delta(Q_{1},\mathrm{W})))$?

### v3

@@ 3 The quantifier $Q_{1}$
As a second example, let $Q_{1}$ be the quantifier asserting “there are $\geq\aleph_{1}$ many…”. It is well-known that $\Delta(\mathcal{L}(Q_{1}))\neq\mathcal{L}(Q_{1})$, as demonstrated by the property of an equivalence relation having uncountably many uncountable classes (see [BF85, Chapter II section 7.2]). This raises the question, what is the relationship between $C(\mathcal{L}(Q_{1}))$ and $C(\Delta(\mathcal{L}(Q_{1})))$?
@@
The proof will use the following machinery. Recall that an Aronszajn tree is a tree of height $\omega_{1}$ such that all its levels are countable, and it has no uncountable chain (a branch); a Suslin tree is an Aronszajn tree which has no uncountable antichain; and an Aronszajn tree $S$ is called special if there is a specializing function – a function $c:S\to A$ such that $A$ is countable, and for every $q,p\in S$, if $c(q)=c(p)$ then $q,p$ are incompatible in $S$. Note that a special Aronszajn tree must have an uncountable chain, so cannot be Suslin. In general, there might be Aronszajn trees that are neither special nor Suslin.
A Suslin tree can be considered as a forcing notion which has the countable chain condition (ccc), and adds a generic uncountable chain to itself. Furthermore, if $S,T$ are Suslin trees such that the tree product $S\otimes T$ is Suslin as well, then forcing with one preserves the Suslinity of the other. Note that, since the tree product $S\otimes T$ is dense in the forcing product $S\times T$, this corresponds to the following well-known fact (see [Jec07, Lemma 16.6]):
Fact 17.
[…]
To summarize, we have the following properties:
Lemma 18.
If $T,S$ are Suslin trees such that $T\otimes S$ is Suslin, then
  1. 1.
[…]
  2. 2.
[…]
Suslin trees satisfying the conclusion of the lemma are called independent. By [AS93, Theorem 7.3], there is a sequence of Suslin trees $\mathcal{T}=\left\langle T_{n}\mid n<\omega\right\rangle\in L$ such that for every distinct $i_{1},...,i_{k}\in\omega$, the product tree $T_{i_{1}}\otimes...\otimes T_{i_{k}}$ is Suslin. To summarize, for every $i\in\omega$:
[…]
  • $T_{i},\mathcal{S}(T_{i})\Vdash$ “for every distinct $i_{1},...,i_{k}\in\omega\smallsetminus\{i\}$, the product tree $T_{i_{1}}\otimes...\otimes T_{i_{k}}$ is Suslin”.
So, all in all, we can either force a branch or specialize $T_{i}$ without affecting the Suslinity of any of the other $T_{j}$s or the independence between them. Note that all posets used, as well as their finite products, are ccc, and as a consequence a finite support product of them preserves $\omega_{1}$.
@@
Let $a\subset\omega$ be a Cohen real over $L$. Using a finite support product, we force over $L[a]$ to add a branch to any $T_{i}$ for $i\in a$ (using $T_{i}$ itself) and specialize any $T_{j}$ for $j\notin a$ (using $\mathcal{S}(T_{j})$). As we noted, this product preserves $\omega_{1}$. Then we force Martin’s Axiom ($\mathsf{MA}$) in the standard way (as in [ST71]) over the resulting model, and denote the final model $W$. Note that still, $\omega_{1}^{W}=\omega_{1}^{L}$. We now work in $W$. As a consequence of $\mathsf{MA}$, any Aronszajn tree in $W$ is special ([BMR70]), and we have
@@ 6 Conclusion and open questions
Another set of questions regards the question of idempotence: for a logic $\mathcal{L}$, we say that $C(\mathcal{L})$ is provably idempotent if it is provable that $C(\mathcal{L})^{C(\mathcal{L})}=C(\mathcal{L})$; and consistently idempotent (and $\neq L$) if it is consistent that $C(\mathcal{L})^{C(\mathcal{L})}=C(\mathcal{L})\neq L$. We will usually omit the phrase in parentheses. Note that the last statement is equivalent to “it is consistent that $V=C(\mathcal{L})\neq L$ holds”. Of course it is only interesting in the case that consistently, $C(\mathcal{L})\neq L$, as $L$ is provably idempotent. Other examples of provable idempotence are models constructed from infinitary logics $C(\mathcal{L}_{\kappa,\kappa})$ [Cha71], but this is the exception rather than the rule, as models such as $\operatorname{HOD},C^{*},C(\mathtt{aa})$ are only consistently idempotent (see [McA71, McA74, KMV21, KMV25, Ya’23, Ya’25]).
It is likely that the constructions we produced here for the logics $C(\Delta(Q_{1}))$ and $C(\Delta(Q_{1},\mathrm{W}))$ result in models where idempotence fails, as, intuitively, we only code the truth of some fact (for example which trees are not Aronszajn), but not the witnesses to this fact (the generic branches). However proving that this is in fact the case requires further research, and we leave the following open:
Question 35.
Are the models $C(\Delta(Q_{1}))$ and $C(\Delta(Q_{1},\mathrm{W}))$ consistently or even provably idempotent?